# Landau-Lifshitz damping from Lindbladian dissipation in quantum magnets

### v2

@@ 1 Introduction
In parallel, a number of approaches exists which compute the Gilbert damping parameter obtaining reasonable values on the basis of time-dependent or non-equilibrium quantum mechanical approaches addressing the baths explicitly [21, 22, 23, 24, 25, 26, 27, 13, 28, 29]. The resulting master equations, however, do not always amount up to Lindbladian dynamics [30]. Still, a quantum mechanical treatment of the baths must lead to Lindbladian dynamics if its preconditions are met: weak-coupling between system and bath, justification of Born-Markov approximation (bath correlations decay faster than the system’s internal dynamics), and the rotating wave approximation [31]. Thus, it must be possible to derive the damping in the LL(G) equations from standard quantum mechanics for open quantum systems in a regime where both descriptions of dissipation are justified. It is the objective of this paper to provide such a systematically controlled derivation. It is not our goal to describe damping beyond the validity of a Lindblad approach nor do we aim at incorporating non-local entanglement to atomistic spin simulations [10, 32, 33, 34] because we rely on local mean-field theory. But we stress that changes of the spin length as computed from the squares of the linear spin expectation values are captured by local mean-field theories.
The Lindblad formalism is a well-established and successful theory for relaxation of quantum systems under certain conditions. The key foundation is to consider an open quantum system coupled to a bath [31] leading to the Lindblad master equation including a so-called dissipator. If the damping of a magnetic systems can be described by the LL(G) equations, which is governed by a single relaxation rate, the latter should be derivable from Lindbladian dynamics. So far, however, Lindblad dynamics and LL(G) dynamics are not linked by a mathematically rigorous derivation, except that for special systems their outcomes are the same [15]. This situation is unsatisfactory since the spins are quantum objects and cannot have their own relaxation independent from the general quantum theory. In this study, we will fill this gap and link Lindbladian dynamics and LL(G) dynamics under certain mild conditions. We show that the latter results from the former in a general limit based on a local mean-field theory, i.e., without non-local correlations [16]. This is justified in higher dimensions. We identify additional terms in the systematic derivation which describe the change of the length of the magnetization, i.e., $|\vec{m}|$.
@@ 2 Brief Recapitulation of Landau-Lifshitz and Lindblad Formalisms
if the Hamilton function reads $H=-\vec{h}_{0}\cdot\hat{\vec{m}}$ and $\hat{\vec{m}}$ is the normalized magnetization of a spin and $\vec{h}_{0}$ a fixed external magnetic field with $h_{0}=g\mu_{\text{B}}B$. The LLG equation reads
$\frac{d\hat{\vec{m}}}{dt}=\hat{\vec{m}}\times\vec{h}_{0}-\lambda\hat{\vec{m}}\times\frac{d\hat{\vec{m}}}{dt}.$ (1b)
Both equations are equivalent within ${\cal O}(\lambda^{2})$ [35, 12] because (1b) can be transformed to
@@
Inserting (1b) in itself and using the bac-cab rule $a\times(b\times c)=b(a\cdot c)-c(a\cdot b)$ for $a,b,c\in\mathds{R}^{3}$ as well as $\hat{\vec{m}}\cdot\frac{d\hat{\vec{m}}}{dt}=0$ leads to (2). Hence the LL and LLG equation stand for almost the same dynamics except for a renormalization of the time $t\to(1+\lambda^{2})t$. The difference only appears in second order in $\lambda$.
Lindbladian dynamics for an operator $A$ in the Heisenberg picture is given by the adjoint Master equation which reads at zero temperature [31, 36]
@@
where the $\gamma_{l}$ are the decay rates of the various dissipative channels which are defined by the Lindblad operators $B_{l}$. It is implied that $B_{l}$ increments the energy in $H$ by $\hbar\omega_{l}>0$ while $B_{l}^{\dagger}$ decreases the energy by the same amount. At finite temperatures, an additional term occurs which is derived and discussed in App. D. Finite temperature does not change the results qualitatively, but increases relaxation and prevents that the magnetizations attains its maximum values.
@@ 3 Single Spin in Time-Dependent Field in Lindblad Formalism
In a first step, we consider a single spin coupled to a static magnetic field which we assume for simplicity to point into $z$ direction $\vec{h}_{0}=h_{0}(0,0,1)^{\top})$. The minimal relaxation channel is given by $B=S^{-}$ with rate $\gamma$. Equation (3) implies
@@
for the spin $S$; with appropriate definitions for longitudinal and transversal relaxation these are the Bloch equations [37]. Qualitatively, they also hold at finite temperatures, see App. D, where also the absorption process induced by $B^{\dagger}$ contributes leading to terms proportional to the bosonic occupation. The equations (4) are exact for $S=1/2$; for general spin they are valid if the spin state deviates only little from the state in the polarized $z$ direction, i.e., if only the magnetic quantum numbers $S$ and $S-1$ occur. We will discuss the justification of this approximation below. The solutions of (4) read $\langle S^{z}\rangle(t)=(\langle S^{z}\rangle(0)-S)\exp(-2\gamma St)+S$ and $\langle S^{+}\rangle(t)=\exp(-(ih_{0}+\gamma S)t)\langle S^{+}\rangle(0)$. This reflects standard longitudinal relaxation combined with damped transverse precession, i.e., the physically expected result. Yet, this behavior is strikingly different from what follows from Eqs. (1)! Yet this is not the end of the story.
As an important intermediate step we consider the magnetic field $\vec{h}$ to be time dependent. For the Lindbladian approach to be valid, the bath dynamics needs to be faster than the relaxation and the internal dynamics (Born-Markov approximation) [31]. In generic solid state systems, the phonons constitute the bath and their characteristic energy scale is set by the Debye frequency $\omega_{\rm Debye}$. Hence, we assume that $\hbar\omega_{\rm Debye}$ is larger than the magnetic coupling $J$. But it is not necessary to specify the physical origin of the bath except that it displays a rapid internal dynamics which allows us to assume that the bath can adapt to $\vec{h}(t)$ at each instant of time quickly, i.e., the Lindblad operator $B(t)$ adapts to the instantaneous magnetic field. Hence the direction of relaxation is given by $\vec{h}(t)$, i.e., towards aligning the spin with its momentary effective field. In App. A we carry out the systematic rotation of Eqs. (4) and we obtain for $\vec{m}\coloneqq\langle\vec{S}\rangle$
@@ 4 Local Mean-Field Approach and its Numerical Evaluation
Here, the next key element enters. We do not attribute the self-reference to an assumed modification of the Schrödinger equation, but to a local mean-field approximation considering the ensemble of spins. For simplicity, we treat a translational invariant ferromagnet in which each spin is subject to the external field $\vec{h}_{0}=h_{0}(0,0,1)^{\top})$ and to the internal exchange field generated by the surrounding spins interacting with the spin under study so that
$\vec{h}=\vec{h}_{0}+J\vec{m},$ (6)
where $J$ represents the sum over all exchange couplings acting on the considered spin. Note that this approach is easily generalized to arbitrary spin textures defining a local effective magnetic field $\vec{h}_{i}\coloneqq\sum_{j}J_{ij}\vec{m}_{j}$. Eventually, one obtains LL equations for each site. If each spin interacts with a sufficiently large number of other spins or if the spin orientation only changes slowly in space coarse graining can be performed as well to introduce spin densities. Thus, a number of extensions is clearly possible.
@@
Clearly, far in the weak field limit for $J/h_{0}=100$ the agreement is excellent; only a minor shift in the relaxation of the $z$ component is discernible, see panel (a). Lowering the ratio $J/h_{0}$ the deviations increase, but the qualitative behavior still is close to the LL dynamics, see panel (b). Figure 1(c) shows that for significant lower ratios the qualitative (!) behavior changes: the $z$ component essentially displays exponential relaxation as one would expect from (4). These numerical observations corroborate the above ypothesis that the LL dynamics results from a Lindblad equations in the limit of weak external fields $h_{0}$ in comparison to the strong internal effective fields $J$. If this limit is reached by increasing $J$ while keeping $h_{0}$ constant, it is important to increase the relaxation rate $\gamma$ as well such that $\lambda=\gamma/J$ remains constant.
At this point, we emphasize that the use of Eq. (6) for large ratios $J/h_{0}$ implies that the spin orientation is always very close to the direction of the effective magnetic field $\vec{h}$, not to $\vec{h}_{0}$. This justifies the use of the approximation necessary to establish (4) for $S>1/2$. For $S=1/2$, the differential equations (4) are exact anyway.
@@ 5 Analytical Derivation of Generalized Landau-Lifshitz Equation
One could use more numerical data for supporting the above hypothesis. But it is by far more general to back the numerical agreement by an analytical argument which captures the dependencies on all parameter. If the LL dynamics is the weak-field limit of the Lindblad dynamics the LL equation must be derivable by a systematic expansion in the external field $h_{0}$. Expanding the right hand side of (5) in linear order in $h_{0}$ using $\vec{n}=\vec{h}/h$ and (6) yields
@@
The second term in Eq. (7) represents a systematic extension of the LL equations for non-saturated magnetizations still fulfilling all required symmetries. Clearly, it goes beyond a purely classical dynamics because the vector length $|\vec{m}|$ will vary as well; it is a consequence of the local mean-field approach. If, however, we start from an initial magnetization that takes its saturation value this term has no effect anymore and $|\vec{m}|=m_{\infty}=S$ holds for all subsequent times. This implies $C=1$ and that the time derivative on the left hand side equals $\frac{d\hat{\vec{m}}}{dt}$ as well as the precession term equals $\hat{\vec{m}}\times\vec{h}_{0}$. The third term on the right hand side represents the LL damping term if we set $\lambda=\gamma/J$. This analytically supports the hypothesis that the LL dynamics results from the Lindblad dynamics in the limit of weak external fields. It is remarkable that in the case of saturation $|\vec{m}|=S$ the time scale $1/\gamma$ disappears completely and only the time scales $1/h_{0}$ and $J/(\gamma h_{0})$ survive. The fact that for a magnetization almost parallel to the external field the change $\frac{d\vec{m}}{dt}$ becomes very small, even vanishing for parallel orientation, is an essential feature of the LL equation. Here it follows automatically from the Lindblad formalism.
Thus, we eventually derived the LL equation (1a) from Lindbladian dynamics and identified the dimensionless damping parameter $\lambda$ by the ratio of the relaxation rate $\gamma$ to the internal energy scale $J$. The latter should be larger than $\gamma$ in order that the weak-coupling assumption in deriving Lindbladian dynamics hold [31]. Hence we can safely assume $\lambda\leq 1$ so that no unphysical behavior results from the LL equation. We cannot derive the LL equation beyond the realm of Lindbladian dynamics. In addition, we studied larger values of $\lambda$ to see whether the LL equation (1a) or the LLG equation (1b) corresponds better to the Lindbladian dynamics (5), see App. C. In accordance with the analytical derivation of the weak-field limit we find that the LL equation represents the correct limit of weak external fields best. We point out, however, that the LL equation is known to display unphysical behavior for $\lambda>1$ so that restrict ourselves to $\lambda<1$.

### v3

@@ 1 Introduction
In parallel, a number of approaches exists which compute the Gilbert damping parameter obtaining reasonable values on the basis of time-dependent or non-equilibrium quantum mechanical approaches addressing the baths explicitly [21, 22, 23, 24, 25, 26, 27, 13, 28, 29]. The resulting master equations, however, do not always amount up to Lindbladian dynamics [30]. Still, a quantum mechanical treatment of the baths must lead to Lindbladian dynamics if its preconditions are met: justified Born-Markov approximation (bath correlations decay faster than the system’s internal dynamics), and the rotating wave approximation [31]. Thus, it must be possible to derive the damping in the LL(G) equations from standard quantum mechanics for open quantum systems in a regime where both descriptions of dissipation are justified. It is the objective of this paper to provide such a systematically controlled derivation. It is not our goal to describe damping beyond the validity of a Lindblad approach nor do we aim at incorporating non-local entanglement to atomistic spin simulations [10, 32, 33, 34] because we rely on local mean-field theory. But we stress that changes of the spin length as computed from the squares of the linear spin expectation values are captured by local mean-field theories.
The Lindblad formalism is a well-established and successful theory for relaxation of quantum systems under certain conditions. The key foundation is to consider an open quantum system coupled to a bath [31] leading to the Lindblad master equation including a so-called dissipator. If the damping of a magnetic systems can be described by the LL(G) equations, which is governed by a single relaxation rate, the latter should be derivable from Lindbladian dynamics. So far, however, Lindblad dynamics and LL(G) dynamics are not linked by a mathematically rigorous derivation, except that for special systems where their outcomes are the same [15] for small deviations from equilibrium. This situation is unsatisfactory: spins are quantum objects so that the Lindblad approach is applicable. But the phenomenological Landau-Lifshitz approach is also well established. Hence, it suggests itself to ask whether and how these two formalisms are linked. In this study, we try to fill this gap and link Lindbladian dynamics and LL(G) dynamics under certain mild conditions. We show that the latter results from the former in a general limit based on a local mean-field theory, i.e., without non-local correlations [16]. This is justified in higher dimensions. We identify additional terms in the systematic derivation which describe the change of the length of the magnetization, i.e., $|\vec{m}|$.
@@ 2 Brief Recapitulation of Landau-Lifshitz and Lindblad Formalisms
if the Hamilton function reads
$H=-\vec{h}_{0}\cdot\hat{\vec{m}}$ (1b)
and $\hat{\vec{m}}$ is the normalized magnetization of a spin and $\vec{h}_{0}$ a fixed external magnetic field with $h_{0}=g\mu_{\text{B}}B$. The LLG equation reads
$\frac{d\hat{\vec{m}}}{dt}=\hat{\vec{m}}\times\vec{h}_{0}-\lambda\hat{\vec{m}}\times\frac{d\hat{\vec{m}}}{dt}.$ (1c)
Both equations are equivalent within ${\cal O}(\lambda^{2})$ [35, 12] because (1c) can be transformed to
@@
Inserting (1c) in itself and using the bac-cab rule $a\times(b\times c)=b(a\cdot c)-c(a\cdot b)$ for $a,b,c\in\mathds{R}^{3}$ as well as $\hat{\vec{m}}\cdot\frac{d\hat{\vec{m}}}{dt}=0$ leads to (2). Hence the LL and LLG equation stand for almost the same dynamics except for a renormalization of the time $t\to(1+\lambda^{2})t$. The difference only appears in second order in $\lambda$.
Lindbladian dynamics for an operator $A$ in the Heisenberg picture is given by the adjoint quantum Master equation which reads at zero temperature [31, 36]
@@
where $H$ is so far a general Hamiltonian, the $\gamma_{l}$ are the decay rates of the various dissipative channels which are defined by the Lindblad operators $B_{l}$. It is implied that $B_{l}$ increments the energy in $H$ by some energy $\hbar\omega_{l}>0$ while $B_{l}^{\dagger}$ decreases the energy by the same amount. This energy does not appear in (3), bit it is relevant at finite temperatures, see App. D. In the weak-coupling derivation of the Lindblad formalism, one assumes that the rates $\gamma_{l}$ are small relative to the time scales of $H$. But it is shown that the Lindblad structure (3) does not depend on the $\gamma_{l}$ being small [31]. Concretely, we first assume that $H$ is given in Eq. (1b). At finite temperatures, an additional term occurs which is derived and discussed in App. D. Finite temperature does not change the results qualitatively, but increases relaxation and prevents that the magnetizations attains its maximum values.
@@ 3 Single Spin in Time-Dependent Field in Lindblad Formalism
In a first step, we consider a single spin coupled to a static magnetic field which we assume for simplicity to point into $z$ direction $\vec{h}_{0}=h_{0}(0,0,1)^{\top}$ in (1b). Below, we will extend this assumption. The minimal relaxation channel is given by $B=S^{-}$ with rate $\gamma$, i.e., there is only one decay channel $l=1$ so that we can omit this index to lighten the notation. Equation (3) implies
@@
for the spin $S$; with appropriate definitions for longitudinal and transversal relaxation these are the Bloch equations [37]. Qualitatively, they also hold at finite temperatures, see App. D, where also the absorption process induced by $B^{\dagger}$ contributes leading to terms proportional to the bosonic occupation. The equations (4) are exact for $S=1/2$; for general spin they are valid if the spin state deviates only little from the state in the polarized $z$ direction, i.e., if only the magnetic quantum numbers $S$ and $S-1$ occur. We will discuss the justification of this approximation below. The solutions of (4) read $\langle S^{z}\rangle(t)=(\langle S^{z}\rangle(0)-S)\exp(-2\gamma St)+S$ and $\langle S^{+}\rangle(t)=\exp(-(ih_{0}+\gamma S)t)\langle S^{+}\rangle(0)$. This reflects standard longitudinal relaxation combined with damped transverse precession, i.e., the physically expected result. This behavior is strikingly different from what follows from Eqs. (1)! Yet this is not the end of the story.
As an important intermediate step we consider the magnetic field $\vec{h}$ to be time dependent, i.e., we pass in (1b) from $\vec{h}_{0}$ to $\vec{h}(t)$ but stick to the single spin case for clarity. For the Lindbladian approach to be valid, the bath dynamics needs to be faster than the relaxation and the Hamiltonian dynamics (Born-Markov approximation) [31]. In generic solid state systems, the phonons constitute the bath and their characteristic energy scale is set by the Debye frequency $\omega_{\rm Debye}$. Hence, we assume that $\hbar\omega_{\rm Debye}$ is significantly larger than $\gamma$ and $|\vec{h}|$. But it is not necessary to specify the physical origin of the bath except that it displays a rapid internal dynamics. We consider the lack of necessity to know more details about the bath an important advantage of the derivation presented because it renders it general. The rapid internal dynamics allows us to assume that the bath can adapt to $\vec{h}(t)$ at each instant of time quickly, i.e., the Lindblad operator $B(t)$ adapts to the instantaneous magnetic field. Hence the direction of relaxation is given by $\vec{h}(t)$, i.e., towards aligning the spin with its momentary effective field. Of course, the instantaneity appears as a strong assumption and future research should investigate which effects a relaxation of this assumption has.
Another question arises at this point, namely how the bath “knows” about the orientation of $\vec{h}$. For instance, the above quoted phonons do not couple to the magnetic field and thus it seems that they cannot induce relaxation towards parallel alignment with the magnetic field. The answer to this question stems from the fundamental level because the issue of the direction of relaxation poses itself already for the case of a constant field $\vec{h}_{0}$: fundamental thermodynamics, e.g., Boltzmann’s H-theorem, teaches us that the total system consisting of the spin and the bath evolves to states of larger entropy. Concretely, if the bath is at zero temperature as we assume here any excess energy flows from the system to the bath. This orientation of the relaxation is incorporated in the above statement that the Lindblad operator $B=S^{-}$ increases the energy in $H$ by some positive amount. By extension to slowly varying external fields, relaxation due to a coupled bath has the preferred direction that energy flows from the system to the bath at each instant of time.
In App. A we carry out the systematic rotation of Eqs. (4) to a time-dependent reference frame and we obtain for $\vec{m}\coloneqq\langle\vec{S}\rangle$
@@ 4 Local Mean-Field Approach and its Numerical Evaluation
Here, the next key element enters. We do not attribute the self-reference in Eqs. (1), i.e., the double occurrence of $\vec{m}$ and/or its temporal derivative, to an assumed modification of the Schrödinger equation, but to a local mean-field approximation considering the ensemble of spins. This means, we pass to the translationally invariant Hamiltonian
$H=-\sum_{i,j}J_{i,j}\vec{S}_{i}\cdot\vec{S}_{j}$ (6)
with dominating positive couplings so that the fully polarized, ferromagnetic states are the ground states. The sum of the couplings
$J:=\sum_{j}J_{i,j}$ (7)
does not depend on $i$ due to the translational invariance. In local mean-field theory, it is sufficient to consider an effective single-site problem, i.e., a single spin in the effective exchange field $\vec{h}_{\text{eff}}=J\langle\vec{S}_{j}\rangle=J\vec{m}$ of its interaction partners. Then each spin is subjected to the external field $\vec{h}_{0}=h_{0}(0,0,1)^{\top})$ and to this effective exchange field so that the total field reads
$\vec{h}=\vec{h}_{0}+J\vec{m}$ (8)
and we can resort to the effective single-site Hamiltonian (1b) with $\vec{h}_{0}\to\vec{h}(t)$ even though we are considering a lattice model. Note that this approach is easily generalized to arbitrary spin textures defining a local effective magnetic field $\vec{h}_{i}\coloneqq\sum_{j}J_{i,j}\langle\vec{S}_{j}\rangle$. Eventually, one obtains Lindblad equations for each site. If each spin interacts with a sufficiently large number of other spins or if the spin orientation only changes slowly in space coarse graining can be performed as well to introduce spin densities. Thus, a number of extensions is clearly possible.
Figure 1: Solutions (solid lines) of Lindbladian (5) with the mean-field self-consistency (8) for the initial length $S$ of the order parameter $\vec{m}$; the initial direction of the magnetization is tilted by $\pi-0.1$ relative to the $z$ direction. The other couplings are $S=1/2$, $\lambda=0.2$, $J$ is given in the panels, and $\gamma=\lambda J$. The dashed curves represent the solution for the same parameters of the LL equation (1a).
@@
Clearly, far in the weak field limit for $J/h_{0}=100$ the agreement is excellent; only a minor shift in the relaxation of the $z$ component is discernible, see panel (a). Lowering the ratio $J/h_{0}$ the deviations increase, but the qualitative behavior still is close to the LL dynamics, see panel (b). Figure 1(c) shows that for significant lower ratios the qualitative (!) behavior changes: the $z$ component essentially displays exponential relaxation as one would expect from (4). These numerical observations corroborate the above hypothesis that the LL dynamics results from a Lindblad equations in the limit of weak external fields $h_{0}$ in comparison to the strong internal effective fields $J$. If this limit is reached by increasing $J$ while keeping $h_{0}$ constant, it is important to increase the relaxation rate $\gamma$ as well such that $\lambda=\gamma/J$ remains constant.
At this point, we emphasize that the use of Eq. (8) for large ratios $J/h_{0}$ implies that the spin orientation is always very close to the direction of the effective magnetic field $\vec{h}$, not to $\vec{h}_{0}$. This justifies a posteriori the use of the approximation necessary to establish (4) for $S>1/2$, i.e., it is consistent to use this approximation here. For $S=1/2$, the differential equations (4) are exact anyway.
@@ 5 Analytical Derivation of Generalized Landau-Lifshitz Equation
One could use more numerical data for supporting the above hypothesis. But it is by far more general to back the numerical agreement by an analytical argument which captures the dependencies on all parameters. If the LL dynamics is the weak-field limit of the Lindblad dynamics the LL equation must be derivable by a systematic expansion in the external field $h_{0}$. Expanding the right hand side of (5) in linear order in $h_{0}$ using $\vec{n}=\vec{h}/h$ and (8) yields
@@
The second term in Eq. (9) represents a systematic extension of the LL equations for non-saturated magnetizations still fulfilling all required symmetries. Clearly, it goes beyond a purely classical dynamics because the vector length $|\vec{m}|$ will vary as well; it is a consequence of the local mean-field approach. If, however, we start from an initial magnetization that takes its saturation value this term has no effect anymore and $|\vec{m}|=m_{\infty}=S$ holds for all subsequent times. This implies $C=1$ and that the time derivative on the left hand side equals $\frac{d\hat{\vec{m}}}{dt}$ as well as the precession term equals $\hat{\vec{m}}\times\vec{h}_{0}$. The third term on the right hand side represents the LL damping term if we set $\lambda=\gamma/J$ as before. This analytically supports the hypothesis that the LL dynamics results from the Lindblad dynamics in the limit of weak external fields. It is remarkable that in the case of saturation $|\vec{m}|=S$ the time scale $1/\gamma$ disappears completely and only the time scales $1/h_{0}$ and $J/(\gamma h_{0})$ survive. Such behavior is not uncommon for complex models in particular limits, e.g., the rest mass becomes unimportant upon passing from the relativistic to the non-relativistic energy-momentum relation. The fact that for a magnetization almost parallel to the external field the change $\frac{d\vec{m}}{dt}$ becomes very small, even vanishing for parallel orientation, is an essential feature of the LL equation. Here it follows automatically from the Lindblad formalism.
Thus, we eventually derived the LL equation (1a) from Lindbladian dynamics and identified the dimensionless damping parameter $\lambda$ by the ratio of the relaxation rate $\gamma$ to the internal energy scale $J$. If one assumes that the Lindbladian is derived for weak-coupling limit [31], i.e., the coupling between the system and the bath is weak it is natural to assume that $\lambda=\gamma/J$ is small. But, as indicated initially on the Lindblad formalism, this is not mandatory as far as the Lindblad formalism is concerned. Yet, we restrict our reasoning to $\lambda\leq 1$ since otherwise the LL equation displays unphysical behavior results indicating that the combination of large relaxation, weak fields, and local mean-field approximation is not reliable enough. Note that the instability of the LL equation for $\lambda\leq 1$ also applies to the extended equation (9) since it is identical to the LL equation for the saturated case $|\vec{m}|=S$.
In addition, we studied larger values of $\lambda$, but still smaller than unity, to see whether the LL equation (1a) or the LLG equation (1c) corresponds better to the Lindbladian dynamics (5), see App. C. In accordance with the analytical derivation of the weak-field limit we find that the LL equation represents the correct limit of weak external fields best. It its argued that the LLG equation (1c) is better suited since it does not bear the risk of unphysical behavior. It is an intriguing question whether an extension of the derivation presented here allows one to establish a link between the Lindblad formalism and the LLG equation. But this issue is beyond the scope of the present article and thus left to future research.
6 Relaxation of the Length of the Order Parameter